OpenStudy (anonymous):

For the system of linear equations below, use graphing to determine in which quadrant the solution lies. x + y = -2 -5x+6y=-30 Quadrant I Quadrant II Quadrant III Quadrant IV

OpenStudy (anonymous):

(x,y) in 1 (-x,y) in 11 (-x,-y) in111 (x,-y) in iv

OpenStudy (anonymous):

which quadrant wold it be tho

OpenStudy (anonymous):

First, let's graph the two lines. The simplest way to do that is to get it into slope-intercept form.\[\text{Slope-Intercept Form: } y = mx + b\]where m = slope and b = y-intercept. So, let's do the first equation. x + y = -2 Original Equation y = -x - 2 Slope-Intercept form of equation Now let's do the second one. -5x + 6y = -30 Original Equation 6y = 5x - 30 Solving for y y = \(\frac56\)x - 5 Slope-intercept form. Do you know how to graph the equations?

OpenStudy (anonymous):

Once you graph it: